OpenStudy (1479186325):

I am doing percents to change. Can someone help me on this problem? 90 is decreased to 75

OpenStudy (anonymous):

So to find the percent of change, first find the difference between the initial value and the new value: \[90 - 75 = 15\] Then we divide this number by our initial value (90), and multiply it by 100, like so: \[\frac{ 15 }{ 90 } \times 100 = ~?\] Finally, just plug this into a calculator to find out your answer @1479186325